# FT-Shield: A Watermark Against Unauthorized Fine-tuning in Text-to-Image Diffusion Models

### v1

@@ 4.5 Robustness of watermark
The robustness of a watermark refers to its ability to remain recognizable after undergoing various modifications, distortions, or attacks. It is a crucial property of watermark because during the images’ circulation, the watermarks may be distorted by some disturbances, such as JPEG compression. The data offender may also use some methods to remove the watermark. In this subsection, we show that our watermarking method can be robust against multiple types of corruptions when proper augmentation is considered in the training of the watermark detector. In the experiment in Table 5, we consider four types of image corruptions including JPEG compression, Gaussian Noise, Gaussian Blur and Random Crop. To make our watermark robust to those corruptions, we consider using all of these four corruptions as an augmentation in the training of each watermark detector. In Table 5, we show the accuracy of the watermark detectors which are trained with or without augmentation on the corrupted images. From the table we can see that, the performance of the watermark detector on the corrupted images is substantially improved after the augmentation is applied in the training of the detector. After the augmentation, the classifier can achieve an accuracy near 100% against all the corruptions in DreamBooth’s images. Even in the images generated by LoRA, where the classifier performs the worst, the accuracy can still be consistently higher than 84%.
@@
4.6 Ablation studies
Reduced watermark rate. Watermark rate refers to the percentage of a dataset which is protected by the watermarks. In real practice, the data protector may have already released their unmarked images before the development of the watermark technique. Therefore, it is necessary to consider the situation where the watermark rate is not 100%. In this subsection, we demonstrate the effectiveness of FT-Shield when the watermark rate is lower than 100%. The experiments are mainly based on the style transfer task using the paintings by artist Louise Abbema. The results are demonstrated in Table 6. As shown by the results, as the percentage of the watermarked images in the training dataset decreases, the watermark detection rate in the generated images also decreases. This is within expectation because when there are less watermarked images in the protected dataset, it is harder for the watermark to be assimilated by the diffusion model. Nonetheless, our method consistently achieves watermark detection much higher than baselines. With a watermark rate of 80%, it can achieve a detection rate close to 100% across all fine-tuning methods. Even when the watermark rate is reduced to 25%, our method can still achieve detection rates higher than 80% across all the fine-tuning methods except LoRA. Although the watermark detection rate for LoRA’s generated images experienced the most substantial decline, it remains much higher than the two baseline methods.
Table 6: Watermark detection rate (TPR) under different watermark rates
WM Rate DreamBooth Textual Inversion Text-to-image LoRA
Ours GW IM Ours GW IM Ours GW IM Ours GW IM
100 % 99.66% 98.47% 91.5% 96.54% 87.78% 88.28% 98.74% 93.33% 88.89% 97.49% 85.83% 87.83%
80 % 99.58% 97.22% 90.68% 97.75% 84.86% 88.35% 96.92% 80.70% 63.73% 95.00% 78.34% 75.21%
50 % 95.92% 94.45% 54.91% 92.92% 83.20% 37.60% 86.25% 77.50% 48.75% 73.75% 66.11% 64.39%
20 % 86.42% 92.78% 13.81% 88.25% 81.25% 19.30% 83.33% 74.45% 32.11% 63.83% 54.45% 44.76%
Ours ($4/255$) Ours ($2/255$)
[…]
Detector trained without data augmentation. As discussed in 3.3, it is necessary to use the generated data from the fine-tuned diffusion model to augment the dataset used for the watermark detector’s training. Table  7 demonstrates the performance of the classifier if there is no augmentation (based on the style transfer task). The classifier is simply trained on the dataset which contains the clean and watermarked protected images. According to the results demonstrated in Table  7, when there is no augmentation, the watermark detector can successfully detect some watermarked images on the generated set, especially those generated by DreamBooth. However, their performance is much worse than the ones with augmented data. This difference demonstrate the necessity to conduct augmentation when training the watermark detector.
5 Conclusions
In this paper, we proposed a novel watermarking method to protect the intellectual property of images against the fine-tuning techniques of text-to-image diffusion models. Our method incorporates the training loss of the fine-tuning of diffusion models in the training objective of the watermarks. This ensures the early assimilation of watermarks by diffusion models, providing a more robust and reliable watermark framework in the protection of image copyright. Empirical results demonstrates the effectiveness of our method and its superiority over the existing watermarking methods.
@@ Appendix B Additional Details about the fine-tuning methods
  • DreamBooth  (Ruiz et al., 2023): DreamBooth is a fine-tuning method to personalize text-to-image diffusion models. It mainly focus on fine-tuning the unet of the diffusion models’ structure with a prior preservation loss to avoid overfitting and language-drift. Whether to update the text-encoder within in the text-to-image model structure is an open option. In the experiment in this paper, we update both the unet and the text-encoder with learning rate to be 2e-6, batch size to be 1 and maximum fine-tuning stpes to be 800. For style transfer task, we use “[V]” as the unique identifier for the specific style and incorporate “[V]” in the prompts in the sampling process to instruct the model to generated images following this style. Similarly, for object transfer, we use ”sks” as the identifier.
  • Textual Inversion (Gal et al., 2022): Textual Inversion is another text-to-image diffusion model personalization method. It focuses on adding a new token which is connected to a specific style or object to the vocabulary of the text-to-image models. This work by using a few representative images to fine-tune the text embedding of the pipeline’s text-encoder. In our experiment, we set the fine-tuning learning rate to be 5.0e-04, batch size to be 1 and maximum fine-tuning steps to be 1500. We use “[V]” as a placeholder to represent the new concepts that the fine-tuning process learn and also incorporate it in the prompts in the sampling process.
@@
  • LoRA (Hu et al., 2021): LoRA works by adding pairs of rank-decomposition matrices to existing weights of the UNet and only train the newly added weights in the fine-tuning process. It also required a image-caption pair dataset for fine-tuning and and the identifier “[V]” in the cations. In experiments, we set the learning rate to be 5e-06, batch size to be 6 and maximum training steps to be 3000.
Appendix C Prompts Used for Images generation
C.1 Prompts Used for style transfer
@@
Sunflowers in the style of [V] The Cliffs of Etelta in the style of [V]
Windmills on the Flower Field in the style of [V] Boats at rest at petit gennevilliers in the style of [V]
Eese in the creek in the style of [V] Meadow with poplars in the style of [V]
Claude haystack at giverny in the style of [V] Olive tree wood in the moreno garden in the style of [V]
@@
Snow scene in the style of [V] Sunrise in the style of [V]
The artist house in the style of [V] The boat studio in the style of [V]
The bodmer oak fontainebleau in the style of [V] The cabin in the style of [V]
Portrait of a Woman with Low Neckline in the style of [V] The sea at saint adresse in the style of [V]
The seine in the style of [V] The sheltered path in the style of [V]
The summer poppy field in the style of [V] Walk in the meadows in the style of [V]
A dog waiting for the owner in the style of [V] Waterloo bridge in the style of [V]
@@
Moonlit night over a calm sea in the style of [V] Majestic castle overlooking a valley in the style of [V]
An old man playing the violin in the style of [V] Wild horses galloping on the shore in the style of [V]
Cherry blossoms in full bloom in the style of [V] Bustling train station in the 1900s in the style of [V]
An Italian vineyard at midday in the style of [V] Birds taking flight from a tree in the style of [V]
@@
The old lighthouse by the cliff in the style of [V] A goat on grass in the style of [V]
Two butterflys in the style of [V] A growling tiger in the style of [V]
@@
An eggplant on vines in the style of [V] Waterfall in the style of [V]
Alm tree at bordighera in the style of [V] Rough sea in the style of [V]
A woman admiring lotus flowers in the style of [V] Two pandas eating bamboos in the style of [V]
Grapes on a vine in the style of [V] A view of mountain in the style of [V]
A man in a suit with a beard in the style of [V] The side face of a red-haired woman in the style of [V]
@@
C.2 Prompts Used for face object transfer
A photo of [V] laughing heartily A photo of [V] by the beach at sunset
A photo of [V] wearing A vintage hat A photo of [V] with A glass of wine
A photo of [V] with A colorful parrot on the shoulder A photo of [V] holding A vintage camera
A photo of [V] holding A coffee cup A photo of [V] on A boat
A photo of [V] holding A bottle of water A photo of [V] wearing A winter scarf and gloves
A photo of [V] holding A bouquet of flowers A photo of [V] wearing oversized sunglasses
A photo of [V] in A library, surrounded by towering bookshelves A photo of [V] in the jungle
A photo of [V] surrounded by festive balloons A photo of [V] amidst colorful autumn leaves
A photo of [V] in A sunny park A photo of [V] in A classroom
A photo of [V] in her bedroom A photo of [V] with A kitten
A photo of [V] with straight black hair A photo of [V] with ear rings
A photo of [V] with blunt-cut bangs A photo of [V] with upset face
A photo of [V] in the street A photo of [V] in front of A window
A photo of [V] with short hair A photo of [V] in front of A flower field
C.3 Prompts Used for lifeless object transfer
A [V] in the snow
A [V] with a wheat field in the background
A [V] on the beach
A [V] with a tree and autumn leaves in the background
A [V] on a cobblestone street
A [V] with the Eiffel Tower in the background
A [V] on top of pink fabric
A [V] on top of green grass with sunflowers around it
A [V] on top of a wooden floor
A [V] on top of a mirror
A [V] with a city in the background
A [V] on top of a dirt road
A [V] with a mountain in the background
A [V] on top of a white rug
A [V] with a blue house in the background
A red [V]
A cube shaped [V]
A [V] placed beside a window
A girl holding a [V]
A [V] on a desk
A [V] on a chair
A [V] beside a computer
A [V] on the top of a roof
A [V] in a box
A [V] on a bed
A man with a [V]
A [V] on a desk
A [V] on a cliff overlooking the sea
A [V] placed on a bookshelf
A [V] under a tree

### v2

@@ 4.6 Robustness of watermark
The robustness of a watermark refers to its ability to remain recognizable after undergoing various modifications, distortions, or attacks. It is a crucial property of watermark because during the images’ circulation, the watermarks may be distorted by some disturbances, such as JPEG compression. The data offender may also use some methods to remove the watermark. In this subsection, we show that our watermark can be robust against multiple types of corruption when proper augmentation is considered in the training of the watermark detector. In the experiment in Table 6, we consider four types of image corruptions including JPEG compression, Gaussian Noise, Gaussian Blur and Random Crop. To make our watermark robust to those corruptions, we consider using all of these four corruptions as an augmentation in the training of each watermark detector. In Table 6, we show the accuracy of FT-Shield-Specific (average of True Positive Rate and True Negative Rate) which are trained with or without augmentation on the corrupted images. The performance of the watermark detector on the corrupted images is substantially improved after the augmentation is applied during the training of the detector. After the augmentation, the classifier can achieve performance near 100% against all the corruptions in DreamBooth’s images. Even in the images generated by LoRA, where the classifier performs the worst, the accuracy can still be consistently higher than 84%.
@@
5 Conclusion
In this paper, we proposed a novel watermarking method to safeguard images’ IP against the fine-tuning of text-to-image diffusion models. To ensure that the watermark can be efficiently and accurately assimilated by the diffusion model, we proposed an algorithm for watermark generation which incorporates the fine-tuning loss of diffusion models in the training loss of watermark. Meanwhile, we introduce a MoE strategy for the watermark detection to enhance its adaptability to diverse fine-tuning methods. Empirical results demonstrates the effectiveness of our method and its superiority over the existing watermarking methods.
@@ Appendix B Additional Details about the Fine-tuning Methods
  • DreamBooth  Ruiz et al. (2023): DreamBooth is a fine-tuning method to personalize text-to-image diffusion models. It mainly focus on fine-tuning the unet of the diffusion models’ structure with a prior preservation loss to avoid overfitting and language-drift. Whether to update the text-encoder within in the text-to-image model structure is an open option. In the experiment in this paper, we update both the unet and the text-encoder with learning rate to be 2e-6, batch size to be 1 and maximum fine-tuning stpes to be 800. For style transfer task, we use “[V]” as the unique identifier for the specific style and incorporate “[V]” in the prompts in the sampling process to instruct the model to generated images following this style. Similarly, for object transfer, we use ”sks” as the identifier.
  • Textual Inversion Gal et al. (2022): Textual Inversion is another text-to-image diffusion model personalization method. It focuses on adding a new token which is connected to a specific style or object to the vocabulary of the text-to-image models. This work by using a few representative images to fine-tune the text embedding of the pipeline’s text-encoder. In our experiment, we set the fine-tuning learning rate to be 5.0e-04, batch size to be 1 and maximum fine-tuning steps to be 1500. We use “[V]” as a placeholder to represent the new concepts that the fine-tuning process learn and also incorporate it in the prompts in the sampling process.
@@
  • LoRA Hu et al. (2021): LoRA works by adding pairs of rank-decomposition matrices to existing weights of the UNet and only train the newly added weights in the fine-tuning process. It also required a image-caption pair dataset for fine-tuning and and the identifier “[V]” in the cations. In experiments, we set the learning rate to 5e-06, batch size to 6 and maximum training steps to 3000.
Appendix C Additional Visualization of Watermarks
In this section, we provide some additional visualization of watermarks as in Figure 5, 6, and 7.
Figure 5: Examples of watermarked images (first line) and images generated through domain adaptation for Pokemon imagery (other lines). The prompt of generation: A robotic cat with wings.
Figure 6: Examples of watermarked images (top row) and images generated from object transfer (clock) shown in subsequent rows. The prompt of generation: A sks clock on top of pink fabric.
Figure 7: Examples of watermarked images (first line) and images generated from face transfer (other lines). The prompt of generation: A photo of sks person wearing a vintage hat.
Appendix D Ablation Studies and Additional Empirical Results
Reduced watermark rate. Watermark rate refers to the percentage of a dataset that is protected by watermarks. In real practice, the data protector may have already released their unmarked images before the development of the watermark technique. Therefore, it is necessary to consider the situation where the watermark rate is not 100%. In this subsection, we demonstrate the effectiveness of FT-Shield when the watermark rate is lower than 100%. The experiments are mainly based on the style transfer task using the paintings by artist Louise Abbema. The results are presented in Table 7. As shown in the table, as the proportion of the watermarked images in the training set decreases, the TPR also decreases. This is within expectation because when there are fewer watermarked images in the protected dataset, it is harder for the watermark to be assimilated by the diffusion model. Nonetheless, our method consistently achieves better performance than baselines. With a watermark rate of 80%, it achieves a detection rate close to 100% across all fine-tuning methods. Even when the watermark rate is reduced to 20%, FT-Shield still maintains detection rates higher than 80% across all the fine-tuning methods except LoRA. Although the watermark detection rate for LoRA’s generated images experienced the most substantial decline, it remains much higher than the two baseline methods.
Table 7: Watermark detection rate (TPR) under different watermark rates (‘FS’ and ‘FM’ refer to ‘FT-Shield-Specific’ and ‘FT-Shield-MoE’, respectively, ‘GW’ stands for ‘Gen-Watermark’ Ma et al. (2023), and ‘DN’ represents ‘DIAGNOSIS’ Wang et al. (2024a)).
WM Rate DreamBooth Textual Inversion Text-to-image LoRA
FS GW FM DN FS GW FM DN FS GW FM DN FS GW FM DN
100 % 99.66% 98.47% 99.67% 91.50% 96.54% 87.78% 96.83% 88.28% 98.74% 93.33% 99.33% 88.89% 97.49% 85.83% 98.33% 87.83%
80 % 99.58% 97.22% 98.83% 90.68% 97.75% 84.86% 95.00% 88.35% 96.92% 80.70% 96.00% 63.73% 95.00% 78.34% 91.67% 75.21%
50 % 95.92% 94.45% 87.33% 54.91% 92.92% 83.20% 88.50% 37.60% 86.25% 77.50% 83.67% 48.75% 73.75% 66.11% 66.17% 64.39%
20 % 86.42% 92.78% 83.50% 13.81% 88.25% 81.25% 85.83% 19.30% 83.33% 74.45% 81.00% 32.11% 63.83% 54.45% 56.50% 44.76%
Ours ($4/255$) Ours ($2/255$)
[…]
Detector trained without data augmentation. As discussed in Section 3, it is necessary to use the generated data from the fine-tuned diffusion model to augment the dataset used for the watermark detector’s training. Table 8 demonstrates the performance of the classifier if there is no augmentation (based on the style transfer task). The classifier is simply trained on the dataset which contains the clean and watermarked protected images. According to the results demonstrated in Table 8, when there is no augmentation, the watermark detector can successfully detect some watermarked images on the generated set, especially those generated by DreamBooth. However, the performance will be much worse than the ones with augmented data. This difference demonstrates the necessity to conduct augmentation when training the watermark detector.
Transferability of tailored watermark detector. We provide some additional results (results for the object transfer tasks) regarding the transferability of tailored watermark detector here. The results are shown in Table 9, indicating that the performance of the detectors is greatly reduced when the classifier is applied to the images generated by a different fine-tuning method.
Robustness of watermark (FT-Shield-MoE). Given the similar performance of FT-Shield-MoE to FT-Shield-Specific, in Section 4.6 we primarily present the detection performance of FT-Shield-Specific. Nonetheless, detection performance of FT-Shield-MoE is also provided, as seen in Table 10. Comparing Table 6 and 10 we can see that, in general, the detection performance of FT-Shield-MoE is similar to that of FT-Shield-Specific. Both FT-Shield-Specific and FT-Shield-MoE achieve good performance in watermark detection on corrupted generated images.
Table 9: Transferability of the watermark detectors in object transfer
budget DreamBooth Textual Inversion
2/255 DreamBooth 98.24% 57.28%
Textual Inversion 68.23% 97.63%
4/255 DreamBooth 98.85% 73.77%
Textual Inversion 72.05% 98.03%
Table 10: Robustness of the watermark against different image corruptions (detected by FT-Shield-MoE)
Corruption DreamBooth Textual Inversion Text-to-image LoRA
Type w/o aug. w/ aug. w/o aug. w/ aug. w/o aug. w/ aug. w/o aug. w/ aug.
JPEG Comp. 62.17% 98.92% 75.42% 98.33% 85.33% 94.33% 76.17% 90.08%
Gaussian Noise 84.58% 98.67% 95.00% 96.67% 71.33% 88.42% 78.33% 85.42%
Gaussian Blur 89.75% 99.17% 69.58% 95.42% 92.17% 94.00% 89.03% 91.50%
Random Crop 82.83% 98.83% 73.75% 95.83% 71.92% 84.08% 73.67% 86.92%
Appendix E Prompts Used for Images Generation
In the following, we provide the prompts used in image generation in our experiments.
E.1 Prompts Used for Face Object Transfer
{adjustwidth}
0cm-1.1cm
A photo of [V] laughing heartily A photo of [V] by the beach at sunset
A photo of [V] wearing a vintage hat A photo of [V] with a glass of wine
A photo of [V] with ear rings A photo of [V] holding a vintage camera
A photo of [V] holding a coffee cup A photo of [V] holding a bouquet of flowers
A photo of [V] in a classroom A photo of [V] wearing A winter scarf and gloves
A photo of [V] on a boat A photo of [V] wearing oversized sunglasses
A photo of [V] in the jungle A photo of [V] in front of a flower field
A photo of [V] with a kitten A photo of [V] amidst colorful autumn leaves
A photo of [V] in a sunny park A photo of [V] holding a bottle of water
A photo of [V] in her bedroom A photo of [V] surrounded by festive balloons
A photo of [V] with upset face A photo of [V] with a colorful parrot on the shoulder
A photo of [V] with blunt-cut bangs A photo of [V] with straight black hair
A photo of [V] in the street A photo of [V] in front of a window
A photo of [V] with short hair
A photo of [V] in a library, surrounded by towering bookshelves
E.2 Prompts Used for Lifeless Object Transfer
{adjustwidth}
0cm-1.1cm
A [V] in the snow A [V] with a wheat field in the background
A [V] on the beach A [V] with a tree and autumn leaves in the background
A [V] on a cobblestone street A [V] with the Eiffel Tower in the background
A [V] on top of pink fabric A [V] on top of green grass with sunflowers around it
A [V] on top of a wooden floor A [V] on top of a mirror
A [V] with a city in the background A [V] on top of a dirt road
A man with a [V] A [V] on top of a white rug
A red [V] A [V] with a blue house in the background
A cube shaped [V] A [V] placed beside a window
A girl holding a [V] A [V] on a desk
A [V] on a chair A [V] beside a computer
A [V] on the top of a roof A [V] in a box
A [V] on a bed A [V] with a mountain in the background
A [V] on a desk A [V] on a cliff overlooking the sea
A [V] placed on a bookshelf A [V] under a tree
E.3 Prompts Used for Style Transfer (Arts)
{adjustwidth}
0cm-1.1cm
@@ E.3 Prompts Used for Style Transfer (Arts)
Sunflowers in the style of [V] A man in a suit with a beard in the style of [V]
Waterfall in the style of [V] Boats at rest at petit gennevilliers in the style of [V]
Eese in the creek in the style of [V] Claude haystack at giverny in the style of [V]
Meadow with poplars in the style of [V] Olive tree wood in the moreno garden in the style of [V]
@@
Snow scene in the style of [V] An Italian vineyard at midday in the style of [V]
The artist house in the style of [V] Cherry blossoms in full bloom in the style of [V]
The cabin in the style of [V] The bodmer oak fontainebleau in the style of [V]
Sunrise in the style of [V] The sea at saint adresse in the style of [V]
The seine in the style of [V] Birds taking flight from a tree in the style of [V]
Walk in the meadows in the style of [V] The summer poppy field in the style of [V]
Rough sea in the style of [V] An old man playing the violin in the style of [V]
@@
A growling tiger in the style of [V] Majestic castle overlooking a valley in the style of [V]
Waterloo bridge in the style of [V] Wild horses galloping on the shore in the style of [V]
The boat studio in the style of [V] Bustling train station in the 1900s in the style of [V]
The sheltered path in the style of [V] Portrait of a Woman with Low Neckline in the style of [V]
@@
A goat on grass in the style of [V] The old lighthouse by the cliff in the style of [V]
Two butterflys in the style of [V] Moonlit night over a calm sea in the style of [V]
@@
An eggplant on vines in the style of [V] Windmills on the Flower Field in the style of [V]
Alm tree at bordighera in the style of [V] A dog waiting for the owner in the style of [V]
A view of mountain in the style of [V] Two pandas eating bamboos in the style of [V]
Grapes on a vine in the style of [V] A woman admiring lotus flowers in the style of [V]
The Cliffs of Etelta in the style of [V] The side face of a red-haired woman in the style of [V]
@@
E.4 Prompts Used for Style Transfer (Pokemon)
{adjustwidth}
0cm-1.1cm A robotic cat with wings A blue and white dinosaur with wings A phoenix with icy feathers A cartoon sunflower with a happy face A steel-spined hedgehog A deer with colorful feathers on it’s head A small, furry creature with large eyes A gray cartoon character with a black tail A fire-breathing fox A pokemon ball with a butterfly on top of it A jellyfish-like creature A red and white toy with a blinking green eye A robotic unicorn with a laser horn A small, insect-like creature with petal-like wings A robotic dog with butterfly wings A magical gnome that walks at night A tree-like creature with glowing eyes A bioluminescent jellyfish with a galaxy pattern inside A creature made of clouds A cartoon character with claws A creature with butterfly-like wings A yellow and orange pokemon with big red eyes A fairy-tale dragon A miniature dragon made of living crystal An armored turtle with a spiked tail A will-o’-the-wisp in the form of a playful kitten A cyborg rabbit with blue eyes A floating island turtle with a mini ecosystem on its back A robotic owl with holographic wings A dog with a wheel in his hand A cyborg penguin with jet engines A mechanical elephant with hover discs for feet A robot bear with solar panels as fur A spectral wolf with aurora-like fur A blue and yellow insect An origami crane that comes to life A pink dog with red ears sitting down A chameleon with digital camouflage A goat with a bell on its head A peacock with holographic tail feathers A red and white dragonfly A saber-toothed cat with plasma claws A blue spider with a green cap A clockwork bird with wings of stained glass A green and yellow bear A starry narwhal with a radiant horn A green and red dragon An ice-cream shaped panda with flavors as fur colors A red and orange pokemon A shadow fox that morphs into smoke A cute animal with horns A wind-up mouse with antique clockwork gears A colorful butterfly A coral reef mermaid with sea glass and shells A pair of owls with orange wings A robot with a green body, yellow arms, and a red head A bamboo dragon that grows real leaves A glowing, ethereal deer with constellation patterns in its fur A neon-furred squirrel with jetpack wings, channeling the futuristic vibe An aquatic creature with the features of a shark and a dolphin